Ofarrell Corporation, a company that produces and sells a single product, has provided its contribution format income statement for March.

  Sales (6,600 units) $429,000
  Variable expenses

297,000

  Contribution margin 132,000
  Fixed expenses

103,500

  Net operating income

$28,500


If the company sells 6,500 units, its net operating income should be closest to:

1.$26,500

2.$28,500

3.27,979

4.$24,000
Answer #1

Contribution per unit is 20.(65-45)

Contribution= 20*6500=130000-103500=26500.

Option 1 is correct